# MG2FlowNet: Accelerating high-Reward Sample Generation via Enhanced MCTS and Greediness Control

### v1

@@ 1 Introduction
Despite these advantages, vanilla GFlowNets often struggle to efficiently discover high reward samples in complex environments. While their inherent exploratory nature enhances diversity, it may lead to excessive coverage of low-reward regions, particularly during early training when the sampling policy lacks guidance and relies on self-collected experience. This results in slow convergence and suboptimal performance in sparse reward scenarios. The fundamental challenge lies in balancing broad exploration with efficient high-reward discovery, highlighting the need for directed exploration strategies that maintain diversity while effectively guiding the model toward high-reward areas. To address these issues, several recent efforts have incorporated reinforcement learning techniques into the GFlowNets framework. Notably, QGFN (Lau et al., 2024) introduces action value (i.e., $Q$-value) to enhance backward policy estimation, while another approach applies MCTS and maximum entropy regularization (Morozov et al., 2024) to enhance planning capabilities. However, these strategies often rely on noisy or inaccurate value estimates in the early training stages and may still fail to effectively guide the model toward high-reward regions. Moreover, they typically lack a fine-grained mechanism for dynamically adjusting the trade-off between exploration and exploitation throughout training. This raises a key question: How can we enhance the model’s ability to explore high-reward regions early in training, while adaptively using historical experience in later stages to maintain high-reward sampling?
@@ 6 Conclusion
We have taken extensive measures to ensure the reproducibility of our results. The main text and appendix provide full details of model architectures, optimization objectives, training hyperparameters, and evaluation metrics. Additional experimental settings, ablation studies, and environment specifications are documented in the supplementary material. We have released the anonymized source code, configuration files, and preprocessing scripts, which are available at https://anonymous.4open.science/r/MG2FlowNet-68B2/. With the released resources and instructions, independent researchers are able to reproduce all reported results reliably.
@@ D.3 Potential Applications of GFlowNets
GFlowNets have demonstrated significant potential across multiple domains due to their unique ability to sample diverse solutions while maintaining reward proportionality. Their strong generalization capabilities enable effective handling of unseen states, making them particularly suitable for exploration-intensive tasks. The technology has shown remarkable success in molecular design, where it outperforms traditional reinforcement learning methods in exploring chemical space while preserving synthetic feasibility and drug-like properties.
Additionally, GFlowNets have emerged as a powerful framework for recommendation systems, demonstrating particular effectiveness in addressing the critical diversity-quality trade-off. Recent studies have successfully applied GFlowNets to enhance listwise recommendations by maintaining recommendation quality while significantly improving diversity Liu et al. (2023b), as well as optimizing user retention through intelligent exploration strategies Liu et al. (2024). Beyond traditional recommendation tasks, GFlowNets have shown remarkable adaptability for large language model fine-tuning across various domains. Notable applications include diverse text generation for sentence infilling and chain-of-thought reasoning Hu et al. (2023), adversarial prompt generation in red teaming scenarios Lee et al. (2024), and complex puzzle-solving in domains such as BlocksWorld and Game24 Yu et al. (2024). These applications collectively demonstrate GFlowNets’ versatility in handling both recommendation tasks and language model optimization challenges.
@@ F.2 Study on Exploration Term
To illustrate the critical role of the exploration term, we use the Hypergrid experiment as an example. As shown in the ablation study results in Table 5, the model performs significantly worse when the exploration term is removed, with performance reduced to less than half of that achieved with the term included. This clearly demonstrates the importance of the exploration component.
@@ F.3 Different Exploration Coefficient $c_{puct}$
As reported in Table 6, empirical results reveal that reducing the exploration coefficient $c_{\text{puct}}$ significantly impairs the model’s ability to consistently generate high reward samples. This observation supports our initial hypothesis: a smaller $c_{\text{puct}}$ limits the model’s capacity to explore less-visited regions of the state space, potentially causing it to overlook promising high-reward areas during the early training phase. On the other hand, setting $c_{\text{puct}}$ to a relatively large value leads to stronger early-stage performance, as it encourages broader exploration and facilitates early discovery of high-reward trajectories. However, as training progresses, such high exploration settings begin to exhibit diminishing returns and even hinder further progress. While early identification of promising regions might be expected to guide subsequent sampling toward them, the problem actually arises from an imbalance between exploration and exploitation. When $c_{\text{puct}}$ is set excessively large, the exploration term in the PUCT (Eq. (5)) selection formula dominates, leading to over-exploration and suboptimal convergence. Based on extensive empirical evaluations, we find that $c_{\text{puct}}=1$ provides a favorable trade-off between exploration and exploitation throughout the training process.
@@ F.4 Comparison of Different Expanding Strategies
In the main text, we mentioned that our expanding strategy is adding all child nodes to the unexpanded node, because the exploration term in our PUCT formulation (Eq. (5)) guarantees that these newly created nodes will be properly prioritized based on their low $n_{visit}$, ensuring they will be systematically explored in future iterations.
@@
Since the strategy of expanding only one child node requires an impractically large number of state visits to discover all 16 modes (rendering it meaningless for comparison), we omit this result here. However, the state visit counts required for discovering 4 and 8 modes clearly demonstrate the infeasibility of single child expansion. Our results show that this approach significantly reduces exploration efficiency.
We attribute this inefficiency to the fundamental limitation of single child expansion: Each training iteration predominantly revisits previously explored nodes due to constrained graph width in MCTS. This severe restriction on new node access dramatically reduces the exploration space. Even in our grid experiment, the state visit counts reached alarming magnitudes, let alone in molecular experiments with exponentially larger state spaces where such costs would become computationally prohibitive.
@@ Appendix H Proof of Equation (10)
By normalizing both numerator and denominator through division by $Q_{max}-Q_{min}$, the formula can be simplified to:
@@ I.1 Limitations
The present study is conducted in controlled environments where both the action space and the reward function remain fixed. This setting is sufficient for validating the core ideas of MG2FlowNet and provides a clear basis for comparison across methods. However, it does not cover scenarios where the available actions evolve during training or where the reward distribution changes over time. These cases are outside the scope of this work and will be investigated in future studies.

### v2

@@ 1 Introduction
Despite these advantages, vanilla GFlowNets often struggle to efficiently discover high-reward samples in complex environments. While their inherent exploratory nature enhances diversity, it may lead to excessive coverage of low-reward regions, particularly during early training when the sampling policy lacks guidance and relies on self-collected experience. This results in slow convergence and suboptimal performance in sparse reward scenarios. The fundamental challenge lies in balancing broad exploration with efficient high-reward discovery, highlighting the need for directed exploration strategies that maintain diversity while effectively guiding the model toward high-reward areas. To address these issues, several recent efforts have incorporated reinforcement learning techniques into the GFlowNets framework. Notably, QGFN (Lau et al., 2024) introduces action value (i.e., $Q$-value) to enhance backward policy estimation, while another approach applies MCTS and maximum entropy regularization (Morozov et al., 2024) to enhance planning capabilities. However, these strategies often rely on noisy or inaccurate value estimates in the early training stages and may still fail to effectively guide the model toward high-reward regions. Moreover, they typically lack a fine-grained mechanism for dynamically adjusting the trade-off between exploration and exploitation throughout training. This raises a key question: How can we enhance the model’s ability to explore high-reward regions early in training, while adaptively using historical experience in later stages to maintain high-reward sampling?
@@ 6 Conclusion
We have taken extensive measures to ensure the reproducibility of our results. The main text and appendix provide full details of model architectures, optimization objectives, training hyperparameters, and evaluation metrics. Additional experimental settings, ablation studies, and environment specifications are documented in the supplementary material. We have released the source code, configuration files, and preprocessing scripts, which are available at https://github.com/ZRNB/MG2FlowNet. With the released resources and instructions, independent researchers are able to reproduce all reported results reliably.
@@ D.3 Potential Applications of GFlowNets
GFlowNets have demonstrated significant potential across multiple domains due to their unique ability to sample diverse solutions while maintaining reward proportionality. Their strong generalization capabilities enable effective handling of unseen states, making them particularly suitable for exploration-intensive tasks. The technology has demonstrated remarkable success in molecular design, outperforming traditional reinforcement learning methods in exploring chemical space while preserving synthetic feasibility and drug-like properties.
Additionally, GFlowNets have emerged as a powerful framework for recommendation systems, demonstrating particular effectiveness in addressing the critical trade-off between diversity and quality. Recent studies have successfully applied GFlowNets to enhance listwise recommendations by maintaining recommendation quality while significantly improving diversity (Liu et al., 2023b), as well as optimizing user retention through intelligent exploration strategies (Liu et al., 2024). Beyond traditional recommendation tasks, GFlowNets have shown remarkable adaptability for fine-tuning large language models across various domains. Notable applications include diverse text generation for sentence infilling and chain-of-thought reasoning (Hu et al., 2023), adversarial prompt generation in red teaming scenarios (Lee et al., 2024), and complex puzzle-solving in domains such as BlocksWorld and Game24 (Yu et al., 2024). These applications collectively demonstrate GFlowNets’ versatility in handling both recommendation tasks and language model optimization challenges.
@@ F.2 Study on Exploration Term
To illustrate the critical role of the exploration term, we use the Hypergrid experiment as an example. As shown in the results of the ablation study in Table 5, the model performs significantly worse when the exploration term is removed, with the performance reduced to less than half of that achieved with the included term. This clearly demonstrates the importance of the exploration component.
@@ F.3 Different Exploration Coefficient $c_{puct}$
As reported in Table 6, empirical results reveal that reducing the exploration coefficient $c_{\text{puct}}$ significantly impairs the model’s ability to consistently generate high reward samples. This observation supports our initial hypothesis: a smaller $c_{\text{puct}}$ limits the model’s capacity to explore less-visited regions of the state space, potentially causing it to overlook promising high-reward areas during the early training phase. On the other hand, setting $c_{\text{puct}}$ at a relatively large value leads to stronger early-stage performance, as it encourages broader exploration and facilitates early discovery of high-reward trajectories. However, as training progresses, such high exploration settings begin to exhibit diminishing returns and even hinder further progress. Although early identification of promising regions might be expected to guide subsequent sampling toward them, the problem actually arises from an imbalance between exploration and exploitation. When $c_{\text{puct}}$ is set excessively large, the exploration term in the PUCT (Eq. (5)) selection formula dominates, leading to over-exploration and suboptimal convergence. Based on extensive empirical evaluations, we find that $c_{\text{puct}}=1$ provides a favorable trade-off between exploration and exploitation throughout the training process.
@@ F.4 Comparison of Different Expanding Strategies
In the main text, we mention that our expansion strategy is to add all child nodes to the unexpanded node, because the exploration term in our PUCT formulation (Eq. (5)) guaranties that these newly created nodes will be properly prioritized based on their low $n_{visit}$, ensuring that they will be systematically explored in future iterations.
@@
Since the strategy of expanding only one child node requires an impractically large number of state visits to discover all 16 modes (rendering it meaningless for comparison), we omit this result here. However, the state visit counts required for discovering 4 and 8 modes clearly demonstrate the infeasibility of single child expansion. Our results show that this approach significantly reduces the efficiency of exploration.
We attribute this inefficiency to the fundamental limitation of single child expansion: Each training iteration predominantly revisits previously explored nodes due to constrained graph width in MCTS. This severe restriction on new node access dramatically reduces the exploration space. Even in our grid experiment, the state visit counts reached alarming magnitudes, let alone in molecular experiments with exponentially larger state spaces, where such costs would become computationally prohibitive.
@@ Appendix H Proof of Equation (10)
By normalizing both the numerator and the denominator through division by $Q_{max}-Q_{min}$, the formula can be simplified to:
@@ I.1 Limitations
The present study is conducted in controlled environments where both the action space and the reward function remain fixed. This setting is sufficient to validate the core ideas of MG2FlowNet and provides a clear basis for comparison across methods. However, it does not cover scenarios where the available actions evolve during training or where the reward distribution changes over time. These cases are outside the scope of this work and will be investigated in future studies.